OpenStudy (anonymous):

Newton's Second Law of Motion is expressed as F = ma. A lawn mower with constant mass is pushed over a length of grass. How much force is required to triple the acceleration rate? A. The force required is one-third the original force. B. The force required is three times the original force. C. The force required remains constant. D. The force required is four times the original.

OpenStudy (anonymous):

@tester97

OpenStudy (tester97):

I dont know that one sorry

OpenStudy (anonymous):

its okay thanks anyway

OpenStudy (anonymous):

@timaashorty

OpenStudy (timaashorty):

Me neither. im never learned this. ;/

OpenStudy (anonymous):

oh thanks anyway! :)

OpenStudy (timaashorty):

Wished I could have helped. I'll tagg @aaronq for you. He's pretty good a t chemistry. Maybe he could help you with is although it's biology.

OpenStudy (anonymous):

thanks @timaashorty

OpenStudy (anonymous):

\[\text{For a constant mass, } \vec{F} \propto \vec{a}\] Acceleration is directly proportional to force, with constant mass. So, if the force were tripled, what do you think should happen to the acceleration?

OpenStudy (timaashorty):

you're welcome [:

OpenStudy (aaronq):

You can rearrange this into \(\dfrac{F}{ma}=0\) Because the mass remains constant we can ignore it, thus we have: \(\dfrac{F_1}{a_1}=\dfrac{F_2}{a_2}\) \(\rightarrow\) \(\dfrac{1}{1}=\dfrac{x}{3}\)

OpenStudy (aaronq):

so x=3, in comparison you need 3 times as much force. They're proportional to one another, given that the mass is constant. ps. post these types of q's in the physics section next time.

OpenStudy (aaronq):

sorry, i didn't see someone was already helping you, i gave it away.

OpenStudy (anonymous):

so the answer is B

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

i already gave you the answer, if you're not going to bother to read what i wrote then i can't help you any further.

OpenStudy (aaronq):

\(\LARGE \color{red}{ in ~comparison ~you ~need~ 3 ~times~ as ~much ~force.}\)

OpenStudy (aaronq):

ps. i don't appreciate you talking smack by attempting to post ridiculous testimonials full of swearing, this is the last time I'm helping you.

OpenStudy (anonymous):

oh snap...